# Quantum Pseudorandomness and Classical Complexity

### v4

@@ 4 Breaking pseudorandomness with a classical oracle
For each $k\in\{0,1\}^{\kappa}$, define $O_{k}=|\varphi_{k}\rangle\langle\varphi_{k}|$. Note that if $X=0$ (i.e. $|\psi\rangle$ is pseudorandom), there always exists a $k$ such that $\mathrm{Tr}(O_{k}|\psi\rangle\langle\psi|)=1$, namely whichever $k$ satisfies $|\psi\rangle=|\varphi_{k}\rangle$. On the other hand, by Lemma 25 and a union bound, if $X=1$ (i.e. $|\psi\rangle$ is Haar-random), $\mathrm{Tr}(O_{k}|\psi\rangle\langle\psi|)<\frac{1}{3}$ for every $k\in\{0,1\}^{\kappa}$, except with probability at most $2^{\kappa}\cdot e^{-(2^{n}-1)/3}\leq\mathrm{negl}(\kappa)$ over $|\psi\rangle$.
If we choose $M=|\{0,1\}^{\kappa}|=2^{\kappa}$, $\varepsilon=\frac{1}{3}$, and $\delta=0.001-2^{k}\cdot e^{-(2^{n}-1)/3}$, then by Theorem 4 there exists a quantum algorithm that takes as input the results $(b,c)$ of $T=O(\kappa)$ single-copy random Clifford measurements, uses the measurement results to estimate $\mathrm{Tr}(O_{k}|\psi\rangle\langle\psi|)$ for each $k$ up to additive error $\frac{1}{3}$, and is correct with probability at least $0.999+2^{\kappa}\cdot e^{-(2^{n}-1)/3}$. In particular, this algorithm can distinguish the pseudorandom ensemble from the Haar-random ensemble, by checking if there exists a $k$ such that the estimate for $\mathrm{Tr}(O_{k}|\psi\rangle\langle\psi|)$ is at least $\frac{2}{3}$. Call this algorithm $\mathcal{C}$, so that $\Pr[\mathcal{C}(b,c)=X]\geq 0.999$.
We will not actually use $\mathcal{C}$, but only its existence. By the optimality of the Bayes decision rule [Ber13, Chapter 4.4.1], because $\mathcal{C}$ uses $(b,c)$ to identify a state $|\psi\rangle$ as either Haar-random or pseudorandom with probability $0.999$, an algorithm that computes the maximum a posteriori estimate of $X$ also succeeds with probability at least $0.999$. In symbols, let $p_{i}=\Pr[X=i\mid b,c]$, which we view as a random variable (depending on $b$ and $c$) for each $i\in\{0,1\}$. Then $\Pr\left[\operatorname*{arg\,max}_{i}p_{i}=X\right]\geq 0.999$.
Next, observe that $\Pr\left[\operatorname*{arg\,max}_{i}p_{i}=X\right]=\mathop{\mathbb{E}}\left[\Pr\left[\operatorname*{arg\,max}_{i}p_{i}=X|b,c\right]\right]=\mathop{\mathbb{E}}\left[\max_{i}p_{i}\right]$, by the law of total expectation. Hence, by Markov’s inequality (and the fact that $\frac{1}{2}\leq\max_{i}p_{i}\leq 1$), we know that $\Pr\left[\max_{i}p_{i}\geq\frac{3}{4}\right]\geq 0.996$. In other words, the Bayes decision rule is usually at least $75\%$ confident in its predictions, so to speak.
Notice that $p_{i}$ equals the probability (conditioned on postselection succeeding) that $\mathcal{B}$ outputs $i$ if it starts with $\rho$ in place of $\hat{\rho}$. For $i\in\{0,1\}$, define $\hat{p}_{i}$ analogously as the postselcted output probabilities of $\mathcal{B}$ itself: $\hat{p}_{i}=\Pr\left[\mathcal{B}(b,c)=i\mid\mathcal{B}(b,c)\in\{0,1\}\right]$. To argue that $\mathcal{A}$ is correct with $0.995$ probability, it suffices to show that
@@ 5 Pseudorandomness from a quantum oracle
In this section, we construct a quantum oracle $(\mathcal{U},\mathcal{P})$ relative to which $\mathsf{BQP}=\mathsf{QMA}$ and PRUs exist.
5.1 $\mathsf{BQP}=\mathsf{QMA}$ relative to $(\mathcal{U},\mathcal{P})$
We start with a lemma showing that the acceptance probability of a quantum query algorithm, viewed as a function of the unitary transformation used in the query, is Lipschitz.
@@ 5.1 $\mathsf{BQP}=\mathsf{QMA}$ relative to $(\mathcal{U},\mathcal{P})$
Let $\mathcal{A}^{U}$ be quantum algorithm that makes $T$ queries to $U\in\mathbb{U}(D)$, and define $f(U)=\Pr\left[\mathcal{A}^{U}=1\right]$. Then $f$ is $2T$-Lipschitz in the Frobenius norm.
@@
Suppose that $||U-V||_{F}\leq d$. By Lemma 8, this implies that the distance between $U$ and $V$ in the diamond norm is at most $2d$. The sub-additivity of the diamond norm under composition implies that as superoperators, $||\mathcal{A}^{U}-\mathcal{A}^{V}||_{\diamond}\leq 2Td$. By 6, it must be the case that $|f(U)-f(V)|\leq 2Td$. ∎
@@
Let $\mathcal{V}^{U}(|\psi\rangle)$ be quantum algorithm that makes $T$ queries to $U\in\mathbb{U}(D)$ and takes as input a quantum state $|\psi\rangle$ on some fixed (but arbitrary) number of qubits. Define $f(U)=\max_{|\psi\rangle}\Pr\left[\mathcal{V}^{U}(|\psi\rangle)=1\right]$. Then $f$ is $2T$-Lipschitz in the Frobenius norm.
@@
$f_{\psi}(U)=\Pr\left[\mathcal{V}^{U}(|\psi\rangle)=1\right],$
so that $f(U)=\max_{|\psi\rangle}f_{\psi}(U)$. Lemma 27 implies that $f_{\psi}$ is $2T$-Lipschitz for every $|\psi\rangle$. Let $U,V\in\mathbb{U}(D)$, and suppose that $|\psi\rangle$ and $|\varphi\rangle$ are such that $f(U)=f_{\psi}(U)$ and $f(V)=f_{\varphi}(V)$. Then:
@@
$\leq 2T||U-V||_{F},$
where the third line uses the fact that $f_{\psi}(V)\leq f_{\varphi}(V)$ and $f_{\varphi}(U)\leq f_{\psi}(U)$, and the last line uses the fact that $f_{\psi}$ and $f_{\varphi}$ are $2T$-Lipschitz. ∎
We are now ready to prove the first main result of this section, that $\mathsf{BQP}^{\mathcal{U},\mathcal{P}}=\mathsf{QMA}^{\mathcal{U},\mathcal{P}}$.
@@
Let $\mathcal{U}=\{\mathcal{U}_{n}\}_{n\in\mathbb{N}}$ be a quantum oracle where each $\mathcal{U}_{n}$ is chosen randomly from $\mu_{2^{n}}^{2^{n}}$. Let $\mathcal{P}$ be an arbitrary $\mathsf{PSPACE}$-complete language. Then with probability $1$ over $\mathcal{U}$, $\mathsf{PromiseBQP}^{\mathcal{U},\mathcal{P}}=\mathsf{PromiseQMA}^{\mathcal{U},\mathcal{P}}$.
@@
First, some notation. We view each $\mathcal{U}_{n}$ alternatively as either a unitary transformation on $2n$ qubits, or as a list of $2^{n}$ different $n$-qubit unitary transformations $\mathcal{U}_{n}=\{\mathcal{U}_{nm}\}_{m\in\{0,1\}^{n}}$ indexed by $n$-bit strings.
@@
Let $\Pi\in\mathsf{PromiseQMA}^{\mathcal{U},\mathcal{P}}$, which means there exists a polynomial-time $\mathsf{PromiseQMA}^{\mathcal{U},\mathcal{P}}$ verifier $\mathcal{V}^{\mathcal{U},\mathcal{P}}(x,|\psi\rangle)$ with completeness $\frac{2}{3}$ and soundness $\frac{1}{3}$. Without loss of generality, we can amplify the completeness and soundness probabilities of $\mathcal{V}$ to $\frac{11}{12}$ and $\frac{1}{12}$, respectively. Let $p(n)$ be a polynomial upper bound on the running time of $\mathcal{V}$ on inputs of length $n$.
@@
We now describe a $\mathsf{PromiseBQP}^{\mathcal{U},\mathcal{P}}$ algorithm $\mathcal{A}^{\mathcal{U},\mathcal{P}}(x)$ such that, with probability $1$ over $\mathcal{U}$, $\mathcal{A}$ computes $\Pi$ on all but finitely many inputs $x\in\mathrm{Dom}(\Pi)$. Let $d=\lfloor\log\left(13824|x|p(|x|)^{2}+2\right)\rfloor$. For each $n\in[d]$, $\mathcal{A}$ performs process tomography on each $\mathcal{U}_{n}$, producing estimates $\widetilde{\mathcal{U}}_{n}$ such that $||\widetilde{\mathcal{U}}_{n}\cdot\widetilde{\mathcal{U}}_{n}^{\dagger}-\mathcal{U}_{n}\cdot\mathcal{U}_{n}^{\dagger}||_{\diamond}\leq\frac{1}{12p(|x|)}$ for every $n$, with probability at least $\frac{2}{3}$ over the randomness of $\mathcal{A}$ [^11] Specifically, one can use the algorithm of [HKOT23] to estimate each $\mathcal{U}_{n}$ to $2^{-\Omega(n)}$ error in diamond norm distance. The estimated unitary transformation $\widetilde{\mathcal{U}}_{n}$ can then be compiled to a circuit using $2^{O(n)}$ $1$- and $2$-qubit gates [VMS04]. Since $n\leq d=O(\log|x|)$, this can be done in polynomial time. Fix $\varepsilon=\frac{1}{12p(|x|)2^{p(|x|)}}$, and let $\mathcal{S}$ be the algorithm from Lemma 19 that samples from an $\varepsilon$-approximate unitary $p(|x|)$-design on $n$ qubits, given as input a random seed $r\sim\{0,1\}^{k_{n}}$ where $k_{n}=\mathrm{poly}(n,|x|)$. Assume further that $\mathcal{S}$ samples from a phase-invariant design (as defined before Lemma 24), which can be accomplished via multiplication by a uniformly random $(p(|x|)+1)$th root of unity.
Consider a $\mathsf{PromiseQMA}^{\mathcal{P}}$ verifier $\widetilde{\mathcal{V}}^{\mathcal{P}}(x,|\psi\rangle)$ that simulates $\mathcal{V}^{\mathcal{U},\mathcal{P}}(x,|\psi\rangle)$ by replacing queries to $\mathcal{U}$ as follows. For each $n\in[p(|x|)]\setminus[d]$, $\widetilde{\mathcal{V}}$ samples a function $f_{n}:\{0,1\}^{n}\to\{0,1\}^{k_{n}}$ from a $2p(|x|)$-wise independent function family. Then, for $n\in[d]$, queries to $\mathcal{U}_{n}$ are replaced by queries to $\widetilde{\mathcal{U}}_{n}$. For $n\in[p(|x|)]\setminus[d]$ and $m\in\{0,1\}^{n}$, queries to $\mathcal{U}_{nm}$ are replaced by queries to $\mathcal{S}(f_{n}(m))$ (i.e. the $m$th unitary in $\mathcal{U}_{n}$ is replaced by an element of the $p(|x|)$-design, selected by $f_{n}(m)$). Consider the $\mathsf{PromiseQMA}^{\mathcal{P}}$ promise problem $\widetilde{\Pi}$ corresponding to $\widetilde{V}$ with completeness $\frac{2}{3}$ and soundness $\frac{1}{3}$. Since $\mathsf{PromiseQMA}^{A}\subseteq\mathsf{PromisePSPACE}^{A}$ for all classical oracles $A$, $\widetilde{\Pi}\in\mathsf{PromisePSPACE}^{\mathcal{P}}=\mathsf{PromisePSPACE}$, so $\mathcal{A}$ can decide $\widetilde{\Pi}(x)$ with a single query to $\mathcal{P}$. $\mathcal{A}$ does this, and outputs $\widetilde{\Pi}(x)$.
@@
We now argue that for any $x$, with high probability over $\mathcal{U}$, $\Pr\left[\mathcal{A}^{\mathcal{U},\mathcal{P}}(x)=\Pi(x)\right]\geq\frac{2}{3}$. It will be convenient to define several hybrid verifiers:
  • $V_{1}=\mathcal{V}$.
  • $V_{2}$: For each $n\in[p(|x|)]\setminus[d]$, chooses a matrix $U_{n}\sim\mu_{2^{n}}^{2^{n}}$. Simulates $V_{1}$, replacing queries to $\mathcal{U}_{n}$ by $U_{n}$ for $n\in[p(|x|)]\setminus[d]$.
  • $V_{3}$: For each $n\in[p(|x|)]\setminus[d]$, samples a function $g_{n}:\{0,1\}^{n}\to\{0,1\}^{k_{n}}$ uniformly at random. Simulates $V_{2}$, replacing queries to $U_{nm}$ by queries to $\mathcal{S}(g_{n}(m))$ for $n\in[p(|x|)]\setminus[d]$ and $m\in\{0,1\}^{n}$.
  • $V_{4}$: For each $n\in[p(|x|)]\setminus[d]$, samples a function $f_{n}:\{0,1\}^{n}\to\{0,1\}^{k_{n}}$ from a $2p(|x|)$-wise independent function family. Simulates $V_{3}$, replacing queries to $g_{n}$ by $f_{n}$.
  • $V_{5}$: Simulates $V_{4}$, replacing queries to $\mathcal{U}_{n}$ by queries to $\widetilde{\mathcal{U}}_{n}$ for $n\in[d]$. Note that $V_{5}$ and $\widetilde{\mathcal{V}}$ are equivalent, and that of these hybrids, $V_{5}$ is the only one whose output depends on the randomness of $\mathcal{A}$ (by way of $\widetilde{\mathcal{U}_{n}}$).
For $i\in[5]$ and a fixed choice of $\mathcal{U}$, define:
$\mathrm{acc}(V_{i})=\max_{|\psi\rangle}\Pr\left[V_{i}(x,|\psi\rangle)=1\right],$
which is well defined by the extreme value theorem. We now bound $|\mathrm{acc}(V_{i})-\mathrm{acc}(V_{i-1})|$ for various $i$:
  • By Lemma 28, because $\mathcal{V}$ makes at most $p(|x|)$ queries, we know that $\mathrm{acc}(V_{1})$, viewed as a function of $(\mathcal{U}_{d+1},\mathcal{U}_{d+2},\ldots,\mathcal{U}_{p(|x|)})$, is $2p(|x|)$-Lipschitz in the Frobenius norm. Hence, by Theorem 9 with $N=13824|x|p(|x|)^{2}+2$, $L=2p(|x|)$, and $t=\frac{1}{12}$, we have that:
@@
    $\Pr_{\mathcal{U}}\left[\left|\mathrm{acc}(V_{1})-\mathrm{acc}(V_{2})\right|\geq\frac{1}{12}\right]$ $\leq 2\exp\left(-\frac{(N-2)t^{2}}{24L^{2}}\right)$
    $=2\exp\left(-\frac{13824|x|p(|x|)^{2}\cdot\frac{1}{144}}{96p(|x|)^{2}}\right)$
    $=2e^{-|x|}.$

### v5

@@ 4 Breaking pseudorandomness with a classical oracle
For each $k\in\{0,1\}^{\kappa}$, define $O_{k}\coloneqq|\varphi_{k}\rangle\langle\varphi_{k}|$. Note that if $X=0$ (i.e. $|\psi\rangle$ is pseudorandom), there always exists a $k$ such that $\mathrm{Tr}(O_{k}|\psi\rangle\langle\psi|)=1$, namely whichever $k$ satisfies $|\psi\rangle=|\varphi_{k}\rangle$. On the other hand, by Lemma 26 and a union bound, if $X=1$ (i.e. $|\psi\rangle$ is Haar-random), $\mathrm{Tr}(O_{k}|\psi\rangle\langle\psi|)<\frac{1}{3}$ for every $k\in\{0,1\}^{\kappa}$, except with probability at most $2^{\kappa}\cdot e^{-(2^{n}-1)/3}$ over $|\psi\rangle$. This probability is negligible in $\kappa$ because $n=\omega(\log\kappa)$, by assumption.
If we choose $M=|\{0,1\}^{\kappa}|=2^{\kappa}$, $\varepsilon=\frac{1}{3}$, and $\delta=0.001-2^{k}\cdot e^{-(2^{n}-1)/3}$, then by Theorem 4 there exists a quantum algorithm that takes as input the results $(b,c)$ of $T=O(\kappa)$ single-copy random Clifford measurements of $|\psi\rangle$, uses the measurement results to estimate $\mathrm{Tr}(O_{k}|\psi\rangle\langle\psi|)$ for each $k$ up to additive error $\frac{1}{3}$, and is correct with probability at least $0.999+2^{\kappa}\cdot e^{-(2^{n}-1)/3}$. In particular, this algorithm can distinguish the pseudorandom ensemble from the Haar-random ensemble, by checking if there exists a $k$ such that the estimate for $\mathrm{Tr}(O_{k}|\psi\rangle\langle\psi|)$ is at least $\frac{2}{3}$. Call this algorithm $\mathcal{C}$, so that $\Pr[\mathcal{C}(b,c)=X]\geq 0.999$.
We will not actually use $\mathcal{C}$, but only its existence. By the optimality of the Bayes decision rule (Lemma 5), because $\mathcal{C}$ uses $(b,c)$ to identify a state $|\psi\rangle$ as either Haar-random or pseudorandom with probability $0.999$, an algorithm that computes the maximum a posteriori estimate of $X$ also succeeds with probability at least $0.999$. In symbols, let $p_{i}=\Pr[X=i\mid b,c]$, which we view as a random variable (depending on $b$ and $c$) for each $i\in\{0,1\}$. Then $\Pr\left[\operatorname*{arg\,max}_{i}p_{i}=X\right]\geq 0.999$.
Next, observe that $\Pr\left[\operatorname*{arg\,max}_{i}p_{i}=X\right]=\mathop{\mathbb{E}}\left[\Pr\left[\operatorname*{arg\,max}_{i}p_{i}=X|b,c\right]\right]=\mathop{\mathbb{E}}\left[\max_{i}p_{i}\right]$, by the law of total expectation. So, by Markov’s inequality (and the fact that $\max_{i}p_{i}\leq 1$), we know $\Pr\left[\max_{i}p_{i}\geq\frac{3}{4}\right]\geq 0.996$. In other words, the Bayes decision rule is usually at least $75\%$ confident in its predictions, so to speak.
Notice that $p_{i}$ equals the probability (conditioned on postselection succeeding) that $\mathcal{B}$ outputs $i$ if it starts with $\rho$ in place of $\hat{\rho}$. For $i\in\{0,1\}$, define $\hat{p}_{i}$ analogously as the postselcted output probabilities of $\mathcal{B}$ itself: $\hat{p}_{i}\coloneqq\Pr\left[\mathcal{B}(b,c)=i\mid\mathcal{B}(b,c)\in\{0,1\}\right]$. To argue that $\mathcal{A}$ is correct with $0.995$ probability, it suffices to show that
@@ 5 Pseudorandomness from a quantum oracle
In this section, we construct a quantum oracle $(\mathcal{U},\mathcal{C})$ relative to which $\mathsf{PromiseBQP}=\mathsf{PromiseQMA}$ and PRUs exist. Let us first describe the oracle, which consists of two parts: a quantum oracle $\mathcal{U}$ and a classical oracle (i.e. a language) $\mathcal{C}$.
5.1 Definition of the oracle
The oracle $\mathcal{U}$ is a sequence of unitaries $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}}$, where each $\mathcal{U}_{n}$ is a direct sum of $2^{n}$ different Haar-random $n$-qubit unitaries. In other words, for each $n$ we sample $\mathcal{U}_{n}\sim\mu_{2^{n}}^{2^{n}}$. We denote this distribution over $\mathcal{U}$ by $\mathcal{U}\sim\mathcal{D}$.
We construct the language $\mathcal{C}$ deterministically and independently of $\mathcal{U}$. We specify the language in stages: first we define $\mathcal{C}$’s behavior on the $1$-bit strings, then the $2$-bit strings, then the $3$-bit strings, and so on. For a string $x$, we define $\mathcal{C}(x)=1$ if the following all hold:
  1. (1)
    $x$ is a description of a quantum oracle circuit $\mathcal{V}^{\overline{\mathcal{U}},\mathcal{C}}(|\psi\rangle)$ that takes a quantum state $|\psi\rangle$ as input, and makes queries to a quantum oracle $\overline{\mathcal{U}}$ and the classical oracle $\mathcal{C}$. Note that $|\psi\rangle$ and $\overline{\mathcal{U}}$ are not part of the description of $\mathcal{V}$; they are auxiliary inputs.
  2. (2)
    $\mathcal{V}$ runs in time at most $|x|-1$, and hence can query $\mathcal{C}$ on inputs of length at most $|x|-1$.
  3. (3)
    The average acceptance probability of $\mathcal{V}$ (viewed as a $\mathsf{QMA}$ verifier) is greater than $1/2$ when averaged over $\overline{\mathcal{U}}\sim\mathcal{D}$. In symbols, we mean precisely:
    $\mathop{\mathbb{E}}_{\overline{\mathcal{U}}\sim\mathcal{D}}\left[\max_{|\psi\rangle}\Pr[\mathcal{V}^{\overline{\mathcal{U}},\mathcal{C}}(|\psi\rangle)=1]\right]>\frac{1}{2}.$
Condition (2) guarantees that $\mathcal{C}$ is not circularly defined, because the quantity in condition (3) depends only on the previously constructed parts of the oracle. Notice also the care we have used in our notation: $\overline{\mathcal{U}}$ is merely used to take an average in the definition of $\mathcal{C}$; it is not the same as $\mathcal{U}$.
5.2 $\mathsf{PromiseBQP}=\mathsf{PromiseQMA}$ relative to $(\mathcal{U},\mathcal{C})$
Now we turn to showing that our oracle satisfies the desired properties. We start with a lemma showing that the acceptance probability of a quantum query algorithm, viewed as a function of the unitary transformation used in the query, is Lipschitz.
@@ 5.2 $\mathsf{PromiseBQP}=\mathsf{PromiseQMA}$ relative to $(\mathcal{U},\mathcal{C})$
Let $\mathcal{A}^{U}$ be a quantum algorithm that makes $T$ queries to $U\in\mathbb{U}(D)$. Define $f:\mathbb{U}(D)\to\mathbb{R}$ by $f(U)\coloneqq\Pr\left[\mathcal{A}^{U}=1\right]$. Then $f$ is $T$-Lipschitz in the Frobenius norm.
@@
Suppose that $||U-V||_{F}\leq d$. Then $||I\oplus U-I\oplus V||_{F}\leq d$, and also $||I\oplus U^{\dagger}-I\oplus V^{\dagger}||_{F}\leq d$, recalling that $I\oplus U$ is controlled-$U$ and $I\oplus V$ is controlled-$V$. By Lemma 9, this implies that the distance between controlled-$U$ and controlled-$V$ in the diamond norm is at most $2d$ (and likewise for controlled-$U^{\dagger}$ and controlled-$V^{\dagger}$). The sub-additivity of the diamond norm under composition implies that as superoperators, $||\mathcal{A}^{U}-\mathcal{A}^{V}||_{\diamond}\leq 2Td$. By 7, we conclude that $|f(U)-f(V)|\leq Td$. ∎
@@
Let $\mathcal{V}^{U}(|\psi\rangle)$ be a quantum algorithm that makes $T$ queries to $U\in\mathbb{U}(D)$ and takes as input a quantum state $|\psi\rangle$ on some fixed (but arbitrary) number of qubits. Define $f:\mathbb{U}(D)\to\mathbb{R}$ by $f(U)\coloneqq\max_{|\psi\rangle}\Pr\left[\mathcal{V}^{U}(|\psi\rangle)=1\right]$. Then $f$ is $T$-Lipschitz in the Frobenius norm.
@@
$f_{\psi}(U)\coloneqq\Pr\left[\mathcal{V}^{U}(|\psi\rangle)=1\right],$
so that $f(U)=\max_{|\psi\rangle}f_{\psi}(U)$. Lemma 28 implies that $f_{\psi}$ is $T$-Lipschitz for every $|\psi\rangle$. Let $U,V\in\mathbb{U}(D)$, and suppose that $|\psi\rangle$ and $|\varphi\rangle$ are such that $f(U)=f_{\psi}(U)$ and $f(V)=f_{\varphi}(V)$. Then:
@@
$\leq T||U-V||_{F},$
where the third line uses the fact that $f_{\psi}(V)\leq f_{\varphi}(V)$ and $f_{\varphi}(U)\leq f_{\psi}(U)$, and the last line uses the fact that $f_{\psi}$ and $f_{\varphi}$ are $T$-Lipschitz. ∎
We are ready to prove the first main result of this section, that $\mathsf{PromiseBQP}^{\mathcal{U},\mathcal{C}}=\mathsf{PromiseQMA}^{\mathcal{U},\mathcal{C}}$.
@@
With probability $1$ over $\mathcal{U}\sim\mathcal{D}$, $\mathsf{PromiseBQP}^{\mathcal{U},\mathcal{C}}=\mathsf{PromiseQMA}^{\mathcal{U},\mathcal{C}}$.
@@
Let $\Pi\in\mathsf{PromiseQMA}^{\mathcal{U},\mathcal{C}}$, which means there exists a polynomial-time verifier $\mathcal{V}^{\mathcal{U},\mathcal{C}}(x,|\psi\rangle)$ with completeness $\frac{2}{3}$ and soundness $\frac{1}{3}$ according to Definition 12. Without loss of generality, we can amplify the completeness and soundness probabilities of $\mathcal{V}$ to $\frac{11}{12}$ and $\frac{1}{12}$, respectively. Let $p(n)$ be a polynomial upper bound on the running time of $\mathcal{V}$ on inputs $x$ of length $n$.
We now describe a $\mathsf{PromiseBQP}^{\mathcal{U},\mathcal{C}}$ algorithm $\mathcal{A}^{\mathcal{U},\mathcal{C}}(x)$ such that, with probability $1$ over $\mathcal{U}$, $\mathcal{A}$ computes $\Pi$ on all but finitely many inputs $x\in\mathrm{Dom}(\Pi)$. The steps of $\mathcal{A}$ are:
  1. (1)
@@
    Let $d\coloneqq\lfloor\log_{2}\left(3456|x|p(|x|)^{2}+2\right)\rfloor$. For each $n\in[d]$, $\mathcal{A}$ performs process tomography on each $\mathcal{U}_{n}$, producing estimates $\widetilde{\mathcal{U}}_{n}$ such that $||\widetilde{\mathcal{U}}_{n}\cdot\widetilde{\mathcal{U}}_{n}^{\dagger}-\mathcal{U}_{n}\cdot\mathcal{U}_{n}^{\dagger}||_{\diamond}\leq\frac{1}{6p(|x|)}$ for every $n$,with probability at least $\frac{2}{3}$ over the randomness of $\mathcal{A}$ [^7] Specifically, one can use the algorithm of [HKOT23] to estimate each $\mathcal{U}_{n}$ to $2^{-\Omega(n)}$ error in diamond norm distance. The estimated unitary transformation $\widetilde{\mathcal{U}}_{n}$ can then be compiled to a circuit using $2^{O(n)}$ $1$- and $2$-qubit gates [VMS04]. Since $n\leq d=O(\log|x|)$, this can be done in polynomial time. Note also that we are using shorthand here: we should really perform process tomography on controlled-$\mathcal{U}_{n}$, so that $||I\oplus\widetilde{\mathcal{U}}_{n}\cdot I\oplus\widetilde{\mathcal{U}}_{n}^{\dagger}-I\oplus\mathcal{U}_{n}\cdot I\oplus\mathcal{U}_{n}^{\dagger}||_{\diamond}\leq\frac{1}{6p(|x|)}$. We use this same shorthand further below. We denote the collection of estimates by $\widetilde{\mathcal{U}}\coloneqq\{\widetilde{\mathcal{U}}_{n}\}_{n\in[d]}$
  2. (2)
    Next, $\mathcal{A}$ constructs a description $x$ of a quantum oracle circuit $\mathcal{W}^{\overline{\mathcal{U}},\mathcal{C}}(x,\widetilde{\mathcal{U}};|\psi\rangle)$. This $\mathcal{W}$ has $x$ and the unitaries in $\widetilde{\mathcal{U}}$ hard-coded into its description, takes an auxiliary input $|\psi\rangle$, [^8] This distinction is why the last argument $|\psi\rangle$ is separated with a semicolon. and queries oracles $\overline{\mathcal{U}}$ and $\mathcal{C}$. On input $|\psi\rangle$, $\mathcal{W}^{\overline{\mathcal{U}},\mathcal{C}}(x,\widetilde{\mathcal{U}};|\psi\rangle)$ replicates the behavior of $\mathcal{V}^{\mathcal{U},\mathcal{C}}(x,|\psi\rangle)$, except that for each $n\in[d]$, queries to $\mathcal{U}_{n}$ are replaced by $\widetilde{\mathcal{U}}_{n}$, and for each $n\in[p(|x|)]\setminus[d]$, queries to $\mathcal{U}_{n}$ are replaced by queries to $\overline{\mathcal{U}}_{n}$.
  3. (3)
    Finally, $\mathcal{A}$ queries $\mathcal{C}(x)$ and outputs the result.
@@
We now show that for any $x\in\mathrm{Dom}(\Pi)$, with high probability over $\mathcal{U}$, $\mathcal{A}$ correctly decides $\Pi$ on $x$, which is to say that $\Pr\left[\mathcal{A}^{\mathcal{U},\mathcal{C}}(x)=\Pi(x)\right]\geq\frac{2}{3}$.
For a fixed $x$, given sequences of unitaries $\widetilde{\mathcal{U}}=\{\widetilde{\mathcal{U}}_{n}\}_{n\in[d]}$ and $\overline{\mathcal{U}}=\{\overline{\mathcal{U}}_{n}\}_{n\in[p(|x|)]\setminus[d]}$, define
$f(\widetilde{\mathcal{U}},\overline{\mathcal{U}})\coloneqq\max_{|\psi\rangle}\Pr\left[\mathcal{W}^{\overline{\mathcal{U}},\mathcal{C}}(x,\widetilde{\mathcal{U}};|\psi\rangle)=1\right].$
Note that, in this notation, $\mathcal{A}$ outputs $1$ if and only if
@@
$\mathop{\mathbb{E}}_{\overline{\mathcal{U}}\sim\mathcal{D}}\left[f(\widetilde{\mathcal{U}},\overline{\mathcal{U}})\right]>\frac{1}{2}.$ (2)
By contrast, the $\mathsf{QMA}$ acceptance probability of $\mathcal{V}$ itself may be written consistently with this notation as:
$f(\mathcal{U},\mathcal{U})=\max_{|\psi\rangle}\Pr\left[\mathcal{V}^{\mathcal{U},\mathcal{C}}(x,|\psi\rangle)=1\right].$ (3)
In effect, our goal is to show that Equation 2 gives a good estimator for Equation 3. We will do so in two steps: we first show that replacing $\mathcal{U}$ in $f$’s second argument with an average over $\overline{\mathcal{U}}$ approximately preserves the $\mathsf{QMA}$ acceptance probability, and then we argue similarly when replacing $\mathcal{U}$ by the estimate $\widetilde{\mathcal{U}}$ in $f$’s first argument.
By Lemma 29, $f$ is $p(|x|)$-Lipschitz with respect to the second argument $\overline{\mathcal{U}}$, viewed as a direct sum of matrices $\overline{\mathcal{U}}\equiv\bigoplus_{n=d+1}^{p(|x|)}\overline{\mathcal{U}}_{n}$. [^9] This is because each query to a single $\overline{\mathcal{U}}_{n}$ may be simulated via one query to the entire direct sum. Hence, from Theorem 10 with $N=3456|x|p(|x|)^{2}+2$, $L=p(|x|)$, and $t=\frac{1}{12}$, we have that:
$\Pr_{\mathcal{U}\sim\mathcal{D}}\left[\left|f(\mathcal{U},\mathcal{U})-\mathop{\mathbb{E}}_{\overline{\mathcal{U}}\sim\mathcal{D}}[f(\mathcal{U},\overline{\mathcal{U}})]\right|\geq\frac{1}{12}\right]$ $\leq 2\exp\left(-\frac{(N-2)t^{2}}{24L^{2}}\right)$
$=2\exp\left(-\frac{3456|x|p(|x|)^{2}\cdot\frac{1}{144}}{24p(|x|)^{2}}\right)$
$=2e^{-|x|}.$ (4)